OpenStudy (anonymous):

I have to use the exponential equation to fill out this graphMonths (x) Population P(x) 0 1 2 3 2 minutes agoDelete 1. Describe how to create the exponential function represented by your table. Use complete sentences and arrive at a final exponential function, P(x). 2. Using your exponential function, P(x), explain the meaning behind each of the numbers and variables in terms of the invasive fish species. 3. The current exponential function P(x) measures the fish population’s growth monthly. Using complete sentences explain how to find the rate of growth every week. 4.

OpenStudy (wolf1728):

Shouldn't there be numbers for the fish population ?

OpenStudy (anonymous):

0 1 1 8 2 64 3 512 thats the #s I have

OpenStudy (anonymous):

forget the (2 minutes agoDelete)

OpenStudy (wolf1728):

The formula is Population = 2^(n*3) n Population 0 1 1 8 2 64 3 512

OpenStudy (anonymous):

yes

OpenStudy (wolf1728):

Here's a graph

OpenStudy (anonymous):

Thanks for the graph but I needed an equation though

OpenStudy (wolf1728):

The equation is in my first posting

OpenStudy (wolf1728):

Population = 2^(n*3)

OpenStudy (ranga):

Exponential growth is of the form y = a * b^x P(x) = a * b^x when x = 0, P = 1 plug it in: 1 = a * b^0 = a * 1 = a a = 1 P(x) = b^x when x = 1, P = 8 8 = b^1 = b b = 8 P(x) = 8^x is the equation.

OpenStudy (ranga):

same as P(x) = 2^(3x) = 8^x

OpenStudy (wolf1728):

Well okay ranga (I just went for the "easy" solution.)

OpenStudy (ranga):

Since they were asking for a description for part 1) I had to go step by step.

OpenStudy (wolf1728):

Yes, I didn't look at the details too much. I just wanted the numbers.

OpenStudy (wolf1728):

Well, I guess it's time to move on.

OpenStudy (ranga):

For part 2) P(x) = 8^x x is the number of months, P(x) is the population of the fish species. The base 8 can also be written as (1 + r) in which case r will be the exponential growth rate. Here P(x) = (1 + 7)^x and therefore r, the growth rate, is 7.

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (ranga):

you are welcome.